# MT-HCCAR: Multi-Task Deep Learning with Hierarchical Classification and Attention-based Regression for Cloud Property Retrieval

### v1

@@ VI Strategies for Selecting the Optimal Model
Based on the evaluation metrics and experiment results expounded in Section V, our analysis extends to the application of the MT models to three different sensors simulated datasets, namely OCI, VIIRS, and ABI. Given the diverse evaluation metrics and the utilization of different satellite datasets, it is imperative to identify the singular optimal model among our multi-task learning-based models, namely MT-CR, MT-HCR, and MT-HCCAR, to be implemented on different datasets.
[…]
@@
While the first and second steps prove effective in selecting the optimal model within the confines of a single metric and a single dataset, extending this evaluation to encompass all datasets and all metrics requires the introduction of absolute performance scores ($P_{ab}$ and 1SE performance scores), denoted as $P_{1SE}$. These proposed scores serve as comprehensive indicators for selecting a single model that excels across diverse datasets and evaluation metrics. Table III shows collected metrics and performance scores for each model and each dataset. The following paragraphs will describe K-fold cross-validation, one standard error rule, and our proposed performance score, respectively.
@@
TABLE III: Model selection using K-fold cross-validation and 1SE rule.
@@
VI-A K-fold cross-validation and 1SE rule
[…]
With the mean values from K-fold cross validation, we then use one standard error rule (1SE rule) [52] to select the most parsimonious model for each dataset. The mean of evaluation metrics, computed over 10 folds, provides an absolute measure of model accuracy. However, it is essential to consider the trade-off between model complexity and absolute performance, as this balance is crucial for assessing the overall quality of a model. Therefore, in addition to the mean value, we also calculate the region defined by one standard error rule (1SE region) of each metric. With the 1SE range of each metric, we need to select “1SE region of the best”, which is the 1SE range of the model with minimum mean error. Then the simplest model whose mean falls within 1SE region of the best is the most parsimonious model according to 1SE rule [53, 43]. Under our experiment setting, the standard error for a certain metric $SE(Metric)$ is calculated as:
$SE(Metric)=\frac{s}{\sqrt{K}}$ (19)
, where $s$ is the square root of variance $s^{2}$ for $Metric$ over 10 folds. Then the 1SE region of each model is defined as: $[\mu-SE,\mu+SE]$.
@@ VI-A K-fold cross-validation and 1SE rule
VI-B The absolute score and the 1SE rule score
@@
$\begin{split}P_{ab}^{\mathcal{M}}&=\sum_{D}\sum_{Metric}P_{ab}^{\mathcal{M,D}}\\ &=\sum_{D}\sum_{Metric}\frac{\mu({Metric})^{\mathcal{M}^{\mathcal{D}}}-\mu({Metric})^{\mathcal{M}^{\mathcal{D}\star}_{\mu}}}{\mu({Metric})^{\mathcal{M}^{\mathcal{D}\star}_{\mu}}}\end{split}$ (20)
, where components $P_{ab}^{\mathcal{M,D}}$ is the score for a specific model $\mathcal{M}$ applied to a specific dataset $\mathcal{D}$, $\mu({Metric})^{\mathcal{M}^{\mathcal{D}}}$ is the mean value of a specific metric $Metric$ of model $\mathcal{M}^{\mathcal{D}}$ applied to dataset $\mathcal{D}$, and $\mathcal{M}^{\mathcal{D}\star}_{\mu}$ is the model with the best mean value for dataset $\mathcal{D}$.
@@
$\begin{split}P_{1SE}^{\mathcal{M}}&=\sum_{D}\mathbf{w_{m}}\sum_{Metric}P_{1SE}^{\mathcal{M,D}}\\ &=\sum_{D}\sum_{Metric}\psi(\mu(Metric),\mu(\mathcal{M},\mathcal{D})\end{split}$ (21)
[…]
VI-C Model selection results
[…]
@@ VI-C Model selection results
VII Conclusions
In this study, we present MT-HCCAR, an end-to-end multi-task learning model tailored for cloud property retrieval on simulated OCI satellite dataset in the PACE project, tackling tasks including cloud masking, cloud phase classification, and Cloud Optical Thickness (COT) prediction. The model is implemented on three sensors’ simulated datasets (OCI, VIIRS, and ABI), respectively, to examine its generalization. Comparative analyses against two baseline methods and ablation studies underscore the effectiveness of the Hierarchical Classification (HC) module and the Classification-Assisted with Attention-based Regression (CAR) module, enhancing performance in both classification and regression tasks. Moreover, our model selection approach, incorporating K-fold cross-validation, the 1SE rule, and two novel performance scores, establishes MT-HCCAR’s superior performance across different datasets and multiple evaluation metrics. The positive results affirm our model’s capability to address real-world challenges in cloud property retrieval and other multi-task applications. Future research endeavors will involve applying the model to actual OCI data, co-located in space and time with VIIRS and ABI post-PACE launch, to assess its performance, and consistency, and enable detailed comparisons with traditional (non-machine-learning) approaches.
Acknowledgment

### v2

@@ Appendix 0.B Cross-Validation and Model Selection
Based on the evaluation metrics and experiment results expounded in Section 5 in the main paper, our analysis extends to the application of the MT models to three different sensors simulated datasets, namely OCI, VIIRS, and ABI. Given the diverse evaluation metrics and the utilization of different satellite datasets, it is imperative to identify the singular optimal model among our multi-task learning-based models, namely MT-CR, MT-HCR, and MT-HCCAR, to be implemented on different datasets.
The process of model selection is based on the four metrics: ACCbi, $AU(\overline{PRC})_{w}$, MSE, and $R^{2}$. The process includes 3 steps: 1) employing K-fold cross-validation, 2) executing one standard error rule (1SE rule), and 3) computing performance scores. Initially, K-fold cross-validation is employed to compare the model’s absolute performance using the mean value of each metric over K folds. This involves aggregating the mean values of all four metrics across models trained on each dataset over K folds. Subsequently, the one standard error rule is applied to assess the balance between model absolute performance and computational complexity.
@@
While the first and second steps prove effective in selecting the optimal model within the confines of a single metric and a single dataset, extending this evaluation to encompass all datasets and all metrics requires the introduction of absolute performance scores ($P_{ab}$ and 1SE performance scores), denoted as $P_{1SE}$. These proposed scores serve as comprehensive indicators for selecting a single model that excels across diverse datasets and evaluation metrics. Table 5 shows collected metrics and performance scores for each model and each dataset. The following paragraphs will describe K-fold cross-validation, one standard error rule, and our proposed performance score, respectively.
@@
Table 5: Model selection using K-fold cross-validation and 1SE rule.
@@
0.B.1 K-fold cross-validation and 1SE rule
K-fold cross-validation is used to compare the absolute performance of each model $\mathcal{M}$ trained by each dataset $\mathcal{D}$ based on mean values of the four evaluation metrics over K folds. We use the number of folds $K=10$ because this value has been found through experiments to generally result in a model skill estimate with low bias and modest variance [1, 2]. For each fold, $(K-1)N/K$ pixels are in the training set, and the left $N/K$ pixels are in the test set. Mean values $\mu(Metric)$ of four evaluation metrics over 10 folds are calculated for each experiment to compare differences in model absolute performance. Among the four metrics, ACCbi and $AU(\overline{PRC})_{w}$ access the classification task while MSE and $R^{2}$ access the regression task. The mean value of a certain metric $Metric$ over 10 folds $\mu(Metric)$ is: $\mu(Metric)=\frac{1}{K}\sum_{k=1}^{K=10}Metric^{k,\mathcal{M}^{\mathcal{D}}}$, where $Metric^{k}$ is $Metric$ for fold $k$ of model $\mathcal{M}$ trained by dataset $\mathcal{D}$.
With the mean values from K-fold cross-validation, we then use one standard error rule (1SE rule) [3] to select the most parsimonious model for each dataset. The mean of evaluation metrics, computed over 10 folds, provides an absolute measure of model accuracy. However, it is essential to consider the trade-off between model complexity and absolute performance, as this balance is crucial for assessing the overall quality of a model. Therefore, in addition to the mean value, we also calculate the region defined by one standard error rule (1SE region) of each metric. With the 1SE range of each metric, we need to select the “1SE region of the best”, which is the 1SE range of the model with minimum mean error. Then the simplest model whose mean falls within the 1SE region of the best is the most parsimonious model according to the 1SE rule [4, 5]. Under our experiment setting, the standard error for a certain metric $SE(Metric)$ is calculated as:
$SE(Metric)=\frac{s}{\sqrt{K}},$ (1)
where $s$ is the square root of variance $s^{2}$ for $Metric$ over 10 folds. Then the 1SE region of each model is defined as: $[\mu-SE,\mu+SE]$.
@@ 0.B.1 K-fold cross-validation and 1SE rule
0.B.2 The absolute score and the 1SE rule score
@@
$\begin{split}P_{ab}^{\mathcal{M}}&=\sum_{D}\sum_{Metric}P_{ab}^{\mathcal{M,D}}\\ &=\sum_{D}\sum_{Metric}\frac{\mu({Metric})^{\mathcal{M}^{\mathcal{D}}}-\mu({Metric})^{\mathcal{M}^{\mathcal{D}\star}_{\mu}}}{\mu({Metric})^{\mathcal{M}^{\mathcal{D}\star}_{\mu}}},\end{split}$ (2)
where components $P_{ab}^{\mathcal{M,D}}$ is the score for a specific model $\mathcal{M}$ applied to a specific dataset $\mathcal{D}$, $\mu({Metric})^{\mathcal{M}^{\mathcal{D}}}$ is the mean value of a specific metric $Metric$ of model $\mathcal{M}^{\mathcal{D}}$ applied to dataset $\mathcal{D}$, and $\mathcal{M}^{\mathcal{D}\star}_{\mu}$ is the model with the best mean value for dataset $\mathcal{D}$.
@@
$\begin{split}P_{1SE}^{\mathcal{M}}&=\sum_{D}\mathbf{w_{m}}\sum_{Metric}P_{1SE}^{\mathcal{M,D}}\\ &=\sum_{D}\sum_{Metric}\psi(\mu(Metric),\mu(\mathcal{M},\mathcal{D}),\end{split}$ (3)
where $\psi(\cdot)$ is a rule that if Model $\mathcal{M}$ trained by dataset $\mathcal{D}$ is selected by 1SE rule based on the mean value of a certain metric, then $\psi(\cdot)$ outputs 1, and otherwise 0. Weights $\mathbf{w_{m}}=[w_{Acc_{bi}},w_{AU(\overline{PRC})_{w}},w_{MSE},w_{R^{2}}]$ define the significance of four metrics. We set $\mathbf{w_{m}}=[1,1,1,1]$ in our experiments because we consider the classification task and regression task equally important in our application. We will further discuss the model performance shown in Table  5 based on 1SE region of the best to show that MT-HCCAR is optimal based on all four metrics of ACCbi, $AU(\overline{PRC})_{w}$, MSE, and $R^{2}$.
0.B.3 Model selection results
This section discusses the superior performance of our proposed MT-HCCAR model, evident in its highest absolute performance score $P_{ab}^{\mathcal{M}}$ score and the top 1SE rule score $P_{1SE}^{\mathcal{M}}$. The last two columns of Table 5 are components of $P_{ab}^{\mathcal{M}}$ and $P_{1SE}^{\mathcal{M}}$. Utilizing Equation 2, the absolute performance scores for each model are calculated as follows: $P_{ab}^{{MT-CR}}=-4.12\%$, $P_{ab}^{{MT-HCR}}=-0.21\%$, and $P_{ab}^{{MT-HCCAR}}=-0.02\%$. The highest $P_{ab}^{\mathcal{M}}$ indicates that MT-HCCAR exhibits the best mean value across all evaluation metrics, including ACCbi, $AU(\overline{PRC})_{w}$, and $R^{2}$, across all satellite datasets.